Explain why (a-b)² = (a+b)² - 4ab
Tell the answer as fast as you could because I'm studying right now.

Grade:10

1 Answers

Sahithi
3 years ago
Given,
(a-b)2 = (a+b)2 – 4ab      (1)
Let’s take R.H.S first and we will  compare it with L.H.S.
(a+b)2 = a2 + b2 +2ab     (2)
By adding  (-4ab) to (2) on both sides we get
(a+b) – 4ab = a2 + b2 +2ab  – 4ab
                      = a2 + b2 – 2ab
                      = (a-b)2
 
Hence prooved